# Effective Two-Stage Knowledge Transfer for Multi-Entity Cross-Domain Recommendation

### v1

@@ 5. CONCLUSION
In this paper, we propose a pre-training & fine-tuning based multi-entity knowledge transfer framework called MKT. MKT can deal with different feature schemas across multiple entities while effectively transferring knowledge from the source entity to the target entity. MKT utilizes a multi-entity pre-training module to extract transferable knowledge across different entities. The heterogeneous feature alignment (HFA) module is designed to scale and align different feature schemas. The polarized distribution loss is used to facilitate an effective extraction of common knowledge. In the end, the extracted common knowledge is adopted for training a target entity model. Experiments conducted on industry datasets demonstrate MKT outperforms state-of-the-art cross-domain recommendation algorithms. MKT has been deployed in Xianyu App, bringing a lift of +4.13% on CTR and +7.07% on UER.

### v2

@@ 5. CONCLUSION
In this paper, we propose a pre-training & fine-tuning based multi-entity knowledge transfer framework called MKT. MKT can deal with different feature schemas across multiple entities while effectively transferring knowledge from the source entity to the target entity. MKT utilizes a multi-entity pre-training module to extract transferable knowledge across different entities. The heterogeneous feature alignment (HFA) module is designed to scale and align different feature schemas. The polarized distribution loss is used to facilitate an effective extraction of common knowledge. In the end, hierarchical transfer of the extracted knowledge is adopted for training a target entity model. Experiments conducted on public and industry datasets demonstrate MKT outperforms state-of-the-art cross-domain recommendation algorithms. MKT has been deployed in Xianyu App, bringing a lift of +4.13% on CTR and +7.07% on UER.